How many gram glucose burn completely to produce sufficient CO2 needed for complete conversion of 12g NaOH into Na2CO3

Grade:11

2 Answers

Vikas TU
6 years ago
Dear Student,
12g of NaOH means 0.3 moles.(since wt of NaOH=40g)
2 moles of NaOH reacts with 1 mole of CO2.
Therefore moles of CO2 required =0.15
1 mole of glucose gives 6 mole of CO2.
moles of glucose required= 0.9
grams of glucose= 0.9*180=162g.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Chintan Shah
5 years ago
 
12g of naoh means 0.3 moles.
1 mole of co2 reacts with 2 moles of naoh.
So,moles of co2 =0.15.
For 1 mole of glucose(c6h12o6) , 6moles of
Co2 . So, for 0.15 moles of co2 , we require 0.025 moles of glucose.
That is 0.025 * 180= 4.5g of glucose. 
  Regards : Chintan T. Shah (vasegdekar academy.)